OpenStudy (anonymous):

The cost of manufacturing x items is given by C(x) = 0.02x^2 + 20x + 1800. Find the marginal cost function. Compuare the marginal cost at x = 2- to the actual cost of producing the 20th item.

OpenStudy (anonymous):

@UnkleRhaukus can u help?

OpenStudy (anonymous):

@amistre64 @Best_Mathematician can u help?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

marginal cost is the derivative of cost

OpenStudy (anonymous):

so i should take the derivative of that?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

actual cost of nth item: C(n) - C(n-1) marginal cost of nth item: C'(n)

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

ait is this derivatives

OpenStudy (amistre64):

Best mistook 2- as something other than a typo ... but was on the right track

OpenStudy (anonymous):

yea its positive 2 sorry

OpenStudy (amistre64):

the marginal cost of the 20th item is, the rate of change in cost of the 20th item actual cost is the extra amount that the 20th item costs compared to the actual cost of the 19th item

OpenStudy (anonymous):

so marginal cost and actual cost r 2 different things

OpenStudy (anonymous):

so after i find the derivative i should plug in 20 into C'(x) to find the marginal cost for 20th items.??

OpenStudy (amistre64):

thats correct, the word marginal in finance relates to derivatives

OpenStudy (amistre64):

yes, C'(20) = marginal cost at x=20

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got it...not a business major but i do have great interest in business

OpenStudy (anonymous):

so that would be my actual cost or?

OpenStudy (anonymous):

for actual cost simply plug in values for marginal cost plugin values in derivatives

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

\[actual~cost=\frac{C(b)-C(a)}{b-a}\] \[marginal~cost=\lim_{a\to b}\frac{C(b)-C(a)}{b-a}\]

OpenStudy (anonymous):

okay i see

OpenStudy (anonymous):

so its asking me for the cost function... For the cost function in problem 2, find the value of x which minimized the average cost

OpenStudy (anonymous):

@amistre64 @Best_Mathematician @electrokid can u help with the second question?

OpenStudy (anonymous):

new question => new post

OpenStudy (anonymous):

ok